Exercise 5-11 (Video)

Pharoah provides shuttle service between four hotels near a medical center and an international airport. Pharoah uses two 10-passenger vans to offer 12 round trips per day. A recent month’s activity in the form of a cost-volume-profit income statement is shown below.
Sales (1,400 passengers) $36,400
Variable costs
    Fuel $4,077
    Tolls and parking 2,507
    Maintenance 696 7,280
Contribution margin 29,120
Fixed costs
    Salaries 11,000
    Depreciation 1,300
    Insurance 4,860 17,160
Net income $11,960
Your answer is incorrect. Try again.
Calculate the break-even point in dollars.
Break-even point $

LINK TO TEXT

LINK TO TEXT

Your answer is incorrect. Try again.
Calculate the break-even point in number of passengers.
Break-even point passengers

LINK TO TEXT

LINK TO TEXT

Your answer is incorrect. Try again.
Without calculations, determine the contribution margin at the break-even point.
Contribution margin at the break-even point $
Click if you would like to Show Work for this question:

Open Show Work
Answer #1

Requirement 1

Break-even point= $21,450

Requirement 2

Break-even point 825 passengers

Requirement 3

Contribution margin at the break-even point $17,160

Working

A Sales $           36,400.00
B Variable Expense $             7,280.00
C=A-B Contribution margin $           29,120.00
D=C/A CM ratio 80%
E Fixed cost $           17,160.00
F=E/D Breakeven sales $           21,450.00
G=36400/1400 Sales price per passenger $                   26.00
H=F/G Break-even point in units                   825

I have used original data for calculating price per passenger.